Renrot

version 1.0, 21 March 2008

Zeus Panchenko

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005, 2006, 2007, 2008 Zeus Panchenko. The Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover texts being “Mailfromd Manual”, and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License”.

(a) The FSF’s Back-Cover Text is: “You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development.”

… Many things we do naturally become difficult only when we try to make them intellectual subjects …

Short Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2 Configuring Renrot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4 Tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 A GNU Free Documentation License. . . . . . . . . . . . . . . . . . . . . . . 25 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

Table of Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Short history of renrot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Typographical conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Acknowledgments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Why RenRot? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Configuring Renrot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2.1 Command Line Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.1 General Command Line Options. . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.2 Aggregation Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.1.3 Contact Sheet Options.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2.2 Configuration File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

3 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

3.1 Template examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.1.1 Filename Template Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.1.2 Aggregation Template Examples. . . . . . . . . . . . . . . . . . . . . . . . . . 15

3.2 Aggregation examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.2.1 Aggregation Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.2.2 Virtual Aggregation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

3.3 Contact Sheet Generation Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.3.1 Contact Sheet Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.3.2 Contact Sheet Ranking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

3.4 Tag Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.4.1 Personal Info Tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.4.2 GPS Tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

4 Tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

Appendix A GNU Free Documentation License

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

A.1 ADDENDUM: How to use this License for your documents. … 32

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

Preface

This edition of the Renrot Manual, last updated 21 March 2008, documents renrot Version 1.0.

Short history of renrot.

Typographical conventions

This manual is written using Texinfo, the GNU documentation formatting language. The same set of Texinfo source files is used to produce both the printed and online versions of the documentation. Because of this, the typographical conventions may be slightly different than in other books you may have read. Examples you would type at the command line are preceded by the common shell primary prompt, ‘$’. The command itself is printed in this font, and the output it produces ‘in this font’, for example:

     $ renrot --version
     RenRot version 1.0

In the text, the command names are printed like this, command line options are displayed in ‘this font’. Some notions are emphasized like this, and if a point needs to be made strongly, it is done this way. The first occurrence of a new term is usually its definition and appears in the same font as the previous occurrence of “definition” in this sentence. File names are indicated like this: ‘/path/to/ourfile’. The variable names are represented like this, keywords and fragments of program text are written in this font.

Acknowledgments

Andy Shevchenko^{1} contributes to the project, makes code design tunning. He is offering and implementing new ideas and helping with Linux related releases.

Phil Harvey (http://www.sno.phy.queensu.ca/~phil/exiftool/), whose

code is the base of renrot EXIF processing code, had offered valued comments and suggestions. ASPLinux company (http://www.asplinux.ru) had included renrot to their distribution. Alex Zasypkin helped with documentation writing and Win32 tests. Fedora Extras Team had included RenRot to its repository. Vyacheslav Garonin proposed idea concerning the virtual aggregation. Sergey Poznyakoff helped with summary and manual style, English and code fixes.

1 coauthor

2 Renrot Anthony Thyssen helped with ImageMagick/PerlMagick related questions. Slaven Rezic had provided invaluable help with perl5.8.x Out of memory ... sbrk() problem on FreeBSD-6.x

1 Description

RenRot is a program to rename and lossless rotate (for now only JPEG format) files according to their EXIF tags values. To prevent incorrect associations, some explanation is needed here. The name of project is short form of ‘REName and ROTate’ and no other interpretation will be used. Renrot is intended to work with a set of files containing EXIF data and can do two things to them: rename and rotate. A set of files can be given either explicitly or using the ‘--extension’ option, which select the files with the given suffix. Renrot operates on files in current working directory, unless given the ‘--work-directory’ option, which changes this default. Renrot renames input files using a flexible name template (which, among others, uses DateTimeOriginal and FileModifyDate EXIF tags, if they exist, otherwise names the file according to the current timestamp). Further, renrot can aggregate files according to the shooting time period or to a given template. Additionally, it rotates files and their thumbnails, as per Orientation EXIF tag. If that tag is absent, the program allows to set rotation parameters using ‘--rotate-angle’ and ‘--rotate-thumb’ command line options. This is currently implemented only for JPEG format. The program can also place commentaries into the following locations:

• Commentary tag from file (see ‘--comment-file’ in see Chapter 2 [Renrot Configuration], page 5)

• UserComment tag from configuration variable (see see Chapter 4 [Tags], page 23)

Personal details may be specified via XMP tags defined in a configuration file, see see Chapter 4 [Tags], page 23. In addition, renrot can aggregate all files in different directories, according to a given date/time pattern template, set with ‘--aggr-template’.

Why RenRot?

Several projects like RenRot are available in the net, but why to choose namely RenRot? Because:

• it does just what it would do - renames and rotates, nothing more than that;

• it is pure CLI with all it’s advantage (no need KDE or any other monster to run);

• it uses Image::ExifTool (the best open tool to work with EXIF data) and libjpeg6 (the best open tool to operate JPEG format files, to correctly rotate both, the entire file and the thumbnail inside it);

• it has very much flex file naming and aggregation template engines; 4 Renrot

• it uses original algorithm of smart Orientation tag rotation;

• it works in batch mode.

2 Configuring Renrot

In the simplest case you will be able to startup renrot without any additional command line options except ‘-e’

2.1 Command Line Options.

2.1.1 General Command Line Options.

‘-c `filename'’
‘--config-file `filename'’

Path to the configuration file.

‘-d directory’
‘--work-directory directory’

Define the working directory.

‘--exclude `filename'’

Specify files to exclude. Wildcards are not allowed. If a set of files is given, there must be as many occurrences of this option as there are files in the set.

‘--sub-fileset `filename'’

Get names of files to operate upon from FILE. The file must contain a file name per line. This option is useful when you need to process only a set of X from Y files in the directory. If specified, the rest of files given in the command line is ignored.

‘-e extention’
‘--extension extention’

Process the files with given extention (JPG, jpeg, CRW, crw, etc). Depending on the operating system, the extension search might or might not be case-sensitive.

‘-n template’
‘--name-template template’

A template to use for creating new file names while renaming. It can also be defined in the configuration file (variable Name Template). The default is ‘%Y%m%d%H%M%S’. For practical uses, see “TEMPLATE EXAMPLES” section. Interpreted sequences are:

%%  a literal %
%#  a literal #

%C Numeric part of the original file name. Implemented for the sake of cameras, that do not supply FileNumber EXIF tag (currently all makes, except *Canon*). Such cameras generate file names starting with letters and ended with digits. No other symbols are allowed in file names, except “-”, “.” and “ ”.

%c Ordinal number of file in the processed file set (see

also ‘--counter-fixed-field’ option).

%d  Day of month (01-31).
%E  The value of ExposureTime tag, if defined.
%e  Old file extension
%F  The value of FNumber tag, if defined.
%H  Hour (00-23).
%I  The value of ISO tag, if defined.

%i FileNumber tag if exists (otherwise, it is replaced by string “NA”).

%M  Minute (00-59).
%m  Month (01-12).

%n Previous filename (the one before renrot started processing).

%O Base part of the original filename (see ‘%o’). In other words, the first part from the beginning to the last dot character.

%o The name file had before it was processed by renrot for the first time. If the file was processed only once, the tag RenRotFileNameOriginal is set to the original file name.

%S  Second (00-59)
%W  The value of WhiteBalance tag, if defined.
%Y  Year with the century (1900, 1901, and so on)

%y  Year without a century (00..99)

You can use value of any EXIF tag to be included as name part. To do that you need to embrace tag name with sign "#", while building name template (see "TEMPLATE EXAMPLES"). Be careful, since any binary EXIF (like ThumbnaiImage) can produce totally unexpected results.

‘--no-rename’

Do not rename files (default is to rename them to YYYYmmddHHMMSS.ext)

‘--counter-fixed-field’
‘--no-counter-fixed-field’

Set fixed length for file counter, used in file name templates (see ‘%c’). It is enabled by default. Use –no-counter-fixed-field to undo its effect.

‘--counter-start number’

Initial value for the file counter (default is 1)

‘--counter-step number’

Step to increment file counter with (default is 1)

‘-r angle’
‘--rotate-angle angle’

Define the angle to rotate files and thumbnails. Allowed values for angle are 90, 180 or 270. It is useful for files not having Orientation tag.

‘--rotate-thumb angle’

Rotate only thumbnails. Allowed values for *ANGLE* are 90, 180 or 270 degrees. Use if the files which were already rotated, but their thumbnails were not.

‘--only-orientation’

Rotate by changing the value of Orientation tag, no real rotation will be made. The sequence of values to rotate an image from normal (0 degrees) by 90 degrees clockwise is: 0 -> 90 -> 180 -> 270 -> 0. It means. set Orientation tag to 90cw after the first rotation, and increase that value by 90 each time the rotation is applied. For 270cw the rotation algorithm uses the reverted sequence. Rotation by 180cw triggers values in two pairs: 0 <-> 180 and 90 <-> 270. This option cannot be applied to mirror values of Orientation tag.

‘--trim’
‘--no-trim’

Pass the "-trim" option to jpegtran(1), to trim if needed. By default, trimming is enabled. Use –no-trim to disable it.

‘--no-rotate’

Do not rotate images (default is to rotate according to EXIF data).

‘--mtime’
‘--no-mtime’

Defines whether to set the file’s mtime, using DateTimeOriginal tag value. Use ‘--no-mtime’ to set it to current time stamp after processing.

‘--no-ren-rot-tag-mtm’
‘--nochg’

Do not rename, rotate, tag and mtime images. It saves files from any changes while allows to do aggregation, contact sheet generation e.t.c.

‘--keywordize’
‘--no-keywordize’

Whether to keywordize. Default is to not. Be careful, since with this option enabled, the existing keywords are rewriten. The keywords are taken from .keywords file or file specified with

option ‘--keywords-file’.
‘-k `filename'’
‘--keywords-file `filename'’

Path to the file with keywords. Its format is a keyword per line. The CR and LF symbols are removed. Empty (only whitespace) lines are ignored. Any leading and trailing whitespace is removed. For example, the line " Test CRLF" is read as " Test ".

‘--keywords-replace’
‘--no-keywords-replace’

Replace existing Keywords tag list rather than add new values to it. Default is not to replace.

‘--comment-file `filename'’

File with commentaries. It is a low priority alias to TagFile =

Comment: `filename'.
‘--user-comment string’
A low priority alias to ‘--tag UserComment: string’
‘-t tag’
‘--tag tag’

See the section "TAGS", for the detailed description

‘--no-tags’

No tags will be written. This is the default.

‘--use-color’
‘--no-use-color’

Colorize output. This does NOT work under Win32.

‘--dry-run’

Do not do anything, only print would have been done.

‘--use-ipc’
‘--no-use-ipc’

Rotate thumbnails using pipe, rather than files. This does NOT work under Win32.

‘-v’

Increase debugging level by 1. Debugging levels from 1 to 4 are internal levels, the levels from 5 till 9 are equivalent to levels 1-5 levels ExifTool with the maximum verbosity for renrot.

‘-?’
‘--help’

Display short usage summary and exit.

‘--version’

Output version information and exit.

2.1.2 Aggregation Options.

‘--aggr-mode mode’

Run aggregation process in given mode. Possible values are: none, delta or template.

‘--aggr-delta number’

Aggregation time delta, in seconds. Files with DateTime-Original and ones of the previous file delta, greater than ‘--aggr-delta’ are placed in the directory, with the name constructed by concatenating the value of the ‘--aggr-directory’ option and the directory name counter.

‘--aggr-directory directory’

Aggregation directory name prefix (default is “Images”)

‘-a template’
‘--aggr-template template’

File name template to use for file aggregation. Images are aggregated by date/time patterns. You may use combination of ‘%d’, ‘%H’, ‘%M’, ‘%m’, ‘%S’, ‘%Y’, and ‘%y’ meta-characters. The template can also be defined in the configuration file (see Aggregation Template variable). The default is ‘%Y%m%d’. For the detailed description, refer to ‘--name-template’ option. For practical uses, see see Section 3.1 [template examples], page 15 section.

‘--aggr-virtual’
‘--no-aggr-virtual’

Defines virtualization for existent aggregation modes. If set, resulting files are placed into the directory given by the command line option ‘--aggr-virtual-directory’ or configuration file option aggregation virtual directory then any changes required by the current aggregation mode are applied. The main effect of ‘--aggr-virtual’ is that any files to be aggregated remain untouched in their places, and symlinks pointing to them are stored in the directory tree created. Use ‘--no-aggr-virtual’ to prevent virtualization.

‘--aggr-virtual-directory directory’

Store virtual aggregation files in directory

2.1.3 Contact Sheet Options.

‘--cs’
‘--contact-sheet’

Create the contact sheet. Currently it works with Thumbnail-Image EXIFs and the files defined as thumbnails (see the option ‘--cs-thm’, below)

‘--cs-file `filename'’
‘--contact-sheet-file `filename'’

Base file name for montage files.

‘--cs-dir directory’
‘--contact-sheet-dir directory’

Temporary directory for montage (created in the begining and deleted at the end of the process)

‘--cs-thm’
‘--contact-sheet-thm’

Files for the montage are already thumbnails

Options bellow are native ImageMagic montage options look ImageMagick documentation for montage options: montage --help

and http://www.imagemagick.org/

Note please, for color^{1} use RGB triplets only like ‘000’ for the ‘black’ or ‘F00’ for the ‘red’.

‘--cs-tile geometry’
‘--contact-sheet-tile geometry’

Tile MxN^{2}.

1 see http://www.imagemagick.org/script/color.php 2 ImageMagick option ‘-tile’

‘--cs-title string’
‘--contact-sheet-title string’

Set the title of the contact sheet^{3}.

‘--cs-bg color’
‘--contact-sheet-bg color’

Background color^{4}.

‘--cs-bd color’
‘--contact-sheet-bd color’

Border color^{5}.

‘--contact-sheet-mt color’

Frame color^{6}.

‘--contact-sheet-fn string’

Render text with this font^{7}.

‘--cs-fl color’
‘--contact-sheet-fl color’

Color to fill the text^{8}.

‘--cs-lb string’
‘--contact-sheet-lb string’

Assign a label to an image^{9}.

‘--cs-fr geometry’
‘--contact-sheet-fr geometry’

Surround image with an ornamental border in N pixels^{10}.

‘--cs-pntsz number’
‘--contact-sheet-pntsz number’

Font point size (IM: ‘-pointsize’).

‘--cs-shadow’
‘--contact-sheet-shadow’

Set the shadow beneath a tile to simulate depth^{11}.

‘--cs-thm-fl color’
‘--contact-sheet-thm-fl color’

Color to fill the text in generated thumbnail.

3 ImageMagick option ‘-title’ 4 ImageMagick option ‘-background’ 5 ImageMagick option ‘-bordercolor’ 6 ImageMagick option ‘-mattecolor’ 7 ImageMagick option ‘-font’ 8 ImageMagick option ‘-fill’ 9 ImageMagick option ‘-label’ 10 ImageMagick option ‘-frame’ 11 ImageMagick option ‘-shadow’

‘--cs-thm-fn string’
‘--contact-sheet-thm-fn string’

Render the generated thumbnail text with this font^{12}.

‘--cs-thm-grfr color’
‘--contact-sheet-thm-grfr color’

Generated thumbnail background gradient color-from

‘--cs-thm-grto color’
‘--contact-sheet-thm-grto color’

Generated thumbnail background gradient color-to

‘--cs-thm-text string’
‘--contact-sheet-thm-text string’

Generated thumbnail text

‘--cs-rank’
‘--contact-sheet-rank’

Run ranking process according the ranks defined with

          ‘--cs-rank-file’

The result is the colored frames of the thumbnails of contact sheets.

‘--cs-rank-file `filename'’
‘--contact-sheet-rank-file `filename'’

Path to the file with ranks. Its format is a "file rankcolor" per line. Filename separated from the color by space or tabulation.

                 01.file.jpg red
                 04.file.JPG green

Only the files found in the file will be ranked.

2.2 Configuration File

A configuration file can be used to set some variables. Renrot looks for its configuration file, named ‘renrot.conf’, in system configuration directories /etc/renrot and </usr/local/etc/renrot>, and in subdirectory .renrot. of the current user home directory. An alternate configuration file can also be explicitly given using the ‘--config-file option’. The configuration file consists of a set of case-insensive keywords and their values separated by equal sign. Each such keyword/value pair occupies a separate line. Boolean variables can have one of the following values: ‘0’, ‘No’, ‘False’, ‘Off’ for ‘false’, and ‘1’, ‘Yes’, ‘True’, ‘On’ for ‘true’. The variables defined for use in configuration file are:

‘use IPC’

Use IPC for execution external command

12 ImageMagick option ‘-font’

‘use color’

Use colorized output (not works under MS Windows) Default color setup for different facilities:

     color = debug: 'green'
     color = error: 'magenta'
     color = fatal: 'red'
     color = info: 'bold'
     color = process: 'white'
     color = warning: 'cyan'
‘mtime’

Set to ‘Yes’ for synchronize mtime with tags, otherwise set it to ‘No’.

‘name template’

File name template (see ‘--name-template’, for the description).

‘trim’

Set to ‘Yes’ to trim rotated images when using jpegtran(1).

‘aggregation mode’

Aggregation mode, possible values are: none, delta or template.

‘aggregation template’

Aggregation template, which defines the file aggregation (see ‘--aggr-template’, for the description).

‘aggregation virtual’

Defines virtualization for the existing aggregation modes (see

the ‘--aggr-virtual’ option).
‘aggregation virtual directory’

Defines a directory for virtual aggregation (see

‘--aggr-virtual-directory’ option).
‘Tag’
‘TagFile’

Refer to the section "TAGS", for the detailed description

‘include’

Include the named file.

Chapter 3: Examples 15

3 Examples

Different aspects examples

3.1 Template examples.

3.1.1 Filename Template Examples

The name template

‘01.%c.%Y%m%d%H%M%S.%i.shtr-#ExposureTime#.f-#FNumber#.iso-#ISO#’

can produce the following names according existent EXIF tags:

     01.01.20030414103656.NA.shtr-1_40.f-2.8.iso-160.jpg
     01.02.20040131230857.100-0078.shtr-1_320.f-2.8.iso-50.jpg
     01.03.20000820222108.NA.shtr-.f-.iso-jpg

3.1.2 Aggregation Template Examples

Look at Aggregation examples

16 Renrot

3.2 Aggregation examples.

3.2.1 Aggregation Examples

The following command:

$ renrot --aggr-mode=template --aggr-template "%Y%m%d" -e jpg

produces the following aggregation: these three files

01.11.20030414103656.NA.jpg
01.12.20030414103813.NA.jpg
01.13.20030414103959.NA.jpg

will be moved to the directory, named ‘20030414’, and

01.14.20040131130857.100-0078.jpg
01.15.20040131131857.100-0079.jpg
01.16.20040131133019.100-0080.jpg

will be moved to the directory, named ‘20040131’.

3.2.2 Virtual Aggregation

If your file system supports links, than you can use virtual aggreagation algorythm. The following command:

$ renrot -e jpg --aggr-mode template --aggr-virtual \
              --aggr-virtual-directory VIRT_AGGR_DIR

produces the following aggregation: for these three files

01.11.20030414103656.NA.jpg
01.12.20030414103813.NA.jpg
01.13.20030414103959.NA.jpg

links will be created in the directory, ‘VIRT_AGGR_DIR/20030414’, and for

01.14.20040131130857.100-0078.jpg
01.15.20040131131857.100-0079.jpg
01.16.20040131133019.100-0080.jpg

links will be created in the directory, named ‘VIRT_AGGR_DIR/20040131’. The very files will remain in the working directory.

Chapter 3: Examples 17

3.3 Contact Sheet Generation Examples

Lets say we have these files we want to make contact sheet for:

$ ls -al
-rw-r--r--  1 zeus  staff  1410248  3 jun  2006 060603.166-6630.JPG
-rw-r--r--  1 zeus  staff  1119839 21 jun  2006 060621.166-6653.JPG
-rw-r--r--  1 zeus  staff  1235431 20 jan  2007 070120.182-8287.JPG
-rw-r--r--  1 zeus  staff  1307705  7 feb  2007 070207.183-8309.JPG
-rw-r--r--  1 zeus  staff  1191954 14 feb  2007 070214.183-8322.JPG
-rw-r--r--  1 zeus  staff   701277  9 jul  2007 070709.192-9265.JPG
-rw-r--r--  1 zeus  staff  1056367 27 jul  2007 070727.192-9291.JPG
-rw-r--r--  1 zeus  staff  1009504 22 aug  2007 070822.193-9342.JPG
-rw-r--r--  1 zeus  staff   702855  1 sen  2007 070901.193-9366.JPG

18 Renrot

3.3.1 Contact Sheet Examples

Simple and almost default variant is:

     $ renrot -e JPG -n '%y%m%d.%i' --cs-title 'CS Sample' --cs-tile '3x3'

the result of the command can be such:

In case we need to make contact sheet file for the images on the file system we have no write access to (like the images written on the CD-ROM), than we have to set contact sheet directory and contact sheet file manually.

     $ renrot -e JPG --cs-dir '/tmp/CS.TMP' --cs-file 'CS.file.jpg'

If in renrot.rc the option ‘contact sheet dir = /tmp/CS.TMP’ was set globally, than you can say only:

     $ renrot -e JPG --cs-file 'CS.file.jpg'

3.3.2 Contact Sheet Ranking

Lets rank thumbnail images in contact sheet with colors

`70120.182-8287.JPG'       OliveDrab2
`070214.183-8322.JPG'      CornflowerBlue
`070727.192-9291.JPG'      plum1
`070901.193-9366.JPG'      plum1
$ renrot -e JPG -n '%y%m%d.%i' --cs-title 'CS Sample' --cs-tile '3x3'
 --cs-rank-file .rank

the result of the command can be such:

Chapter 3: Examples 21

3.4 Tag Examples

3.4.1 Personal Info Tags

Personal information is set to XMP tags. Defaults are undefined.

Tag = CreatorContactInfoCiAdrCity [XMP]: 'Donetsk'
Tag = CreatorContactInfoCiAdrCtry [XMP]: 'UA'
Tag = CreatorContactInfoCiAdrExtadr [XMP]: '1 bld., Some St.'
Tag = CreatorContactInfoCiAdrPcode [XMP]: '12345'
Tag = CreatorContactInfoCiAdrRegion [XMP]: 'Donetsk'
Tag = CreatorContactInfoCiEmailWork [XMP]: 'email at some dot domain'
Tag = CreatorContactInfoCiTelWork [XMP]: '+123456789'
Tag = CreatorContactInfoCiUrlWork [XMP]: 'https://ps.gnu.org.ua/projects/
renrot/'
Tag = Artist [EXIF]: 'John Doe'
Tag = ImageDescription [EXIF]: 'Just to put anything here.'

3.4.2 GPS Tags

For more info on GPS tags consult:

http://www.sno.phy.queensu.ca/~phil/exiftool/TagNames/GPS.html

Following sample is given for Donetsk, Ukraine:

Tag = GPSLatitudeRef [GPS]: 'N'         # N for Noth and S for South
Tag = GPSLatitude [GPS]: '47 59 06'     # in form "%d %.2f %.2f"
Tag = GPSLongitudeRef [GPS]: 'E'        # E for East and W for West
Tag = GPSLongitude [GPS]: '37 46 48'    # in form "%d %.2f %.2f"
Tag = GPSAltitudeRef [GPS]: 'above'     # above as Above, bellow as Bellow
Tag = GPSAltitude [GPS]: '225'          # in form "%d"
Tag = GPSDateStamp [GPS]: ''            # in form "YYYY:MM:DD"
Tag = GPSDateStamp [GPS]: ''            # in form "HH MM SS"

4 Tags

A tag is defined by the following combination: TagName [Group]:'value'. The defined tags are selected to be set and writen to the EXIF tree using the command line option ‘--tag’ and/or configuration file options ‘Tag’. The syntax of the command line option ‘--tag’ is:

     --tag *TagName [Group]: 'value'*

The syntax of the configuration file option ‘Tag’:

     Tag = *TagName [Group]: 'value'*

The parameters ‘TagName’ and ‘Group’ are passed to ExifTool as is. The name of the group must be enclosed in square brackets. Its ‘value’ (after the semicolon) can be enclosed in single quotes. The ‘TagFile’ keyword allows to set multi-line tags from a file. Its syntax is:

     TagFile = *TagName [Group]:* `filename'

The following table summarizes some tags that can be used, with the ‘--tag’ option and ‘Tag’ keyword:

‘Copyright’

Copyright notes.

‘Comment’

General comment.

‘UserComment’

Anything you would like to put as a comment.

‘CreatorContactInfoCiAdrCity’

A city tag.

‘CreatorContactInfoCiAdrCtry’

A country tag.

‘CreatorContactInfoCiAdrExtadr’

Extended address (usually includes street and apartment number).

‘CreatorContactInfoCiAdrPcode’

Zip code.

‘CreatorContactInfoCiAdrRegion’

Region.

‘CreatorContactInfoCiEmailWork’

Email.

‘CreatorContactInfoCiTelWork’

Phone number.

‘CreatorContactInfoCiUrlWork’

URL.

24 Renrot Additionally, you can add any known tag here, using ‘Tag’ or ‘TagFile’ options as described above.

Appendix A GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 59 Temple Place, Suite 330, Boston, MA 02111-1307, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct

from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified

Version, as the publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adja-

cent to the other copyright notices. F. Include, immediately after the copyright notices, a license notice

giving the public permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections

and required Cover Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add

to it an item stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included in the Modified Version.

N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in title with any Invariant Section.

O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties— for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one.

The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. See

http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

A.1 ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

       Copyright (C)  year  your name.
       Permission is granted to copy, distribute and/or modify this document
       under the terms of the GNU Free Documentation License, Version 1.2
       or any later version published by the Free Software Foundation;
       with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
       Texts.  A copy of the license is included in the section entitled ``GNU
       Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

         with the Invariant Sections being list their titles, with
         the Front-Cover Texts being list, and with the Back-Cover Texts
         being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation. If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index 33

Concept Index

This is a general index of all issues discussed in this manual

A

Alex Zasypkin. . . . . . . . . . . . . . . . . . . . . . . . . . 1 Andy Shevchenko . . . . . . . . . . . . . . . . . . . . . . 1 Anthony Thyssen . . . . . . . . . . . . . . . . . . . . . . 1 ASPLinux company . . . . . . . . . . . . . . . . . . . . 1

F

FDL, GNU Free Documentation License

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Fedora Extras Team. . . . . . . . . . . . . . . . . . . . 1

G

GPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

P

Phil Harvey . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

S

Sergey Poznyakoff . . . . . . . . . . . . . . . . . . . . . . 1 Slaven Rezic . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

T

Texinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

V

Vyacheslav Garonin . . . . . . . . . . . . . . . . . . . . 1

X

XPM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21